# Tighter Lower Bounds on Aperiodic Ambiguity Function and Their Asymptotic Achievability

### v2

@@ III-A1 Weight vector A ($\mathbf{w}^{A,q}$)
Corollary 2
For any LAZ $\bm{\Gamma}$ and any integer $q\in[1,Z_{x}]$, we have
${\theta}_{c}^{2}3qZ_{y}(M-1)+{\theta}_{a}^{2}\left(3qZ_{y}-3Z_{y}\right)\geq 3qMNZ_{y}-\left(q^{2}-1\right)MZ_{y}-3N^{2},$ (31a)
${\theta}_{\text{max}}^{2}\geq\frac{3qMNZ_{y}-3N^{2}-q^{2}MZ_{y}+MZ_{y}}{3(qM-1)Z_{y}}.$ (31b)
Proof:
By substituting $\mathbf{w}^{A,q}$ and $D=0$ into (25a),
${\theta}_{c}^{2}(M-1)+{\theta}_{a}^{2}\left(1-\frac{1}{q}\right)$ $\geq M\left(N-\sum_{s,t=0}^{Z_{x}-1}l_{s,t,N}w_{s}w_{t}-\frac{N^{2}}{MZ_{y}}\sum_{i=0}^{Z_{x}-1}w_{i}^{2}\right)$
$=M\left(N-\sum_{s,t=0}^{q-1}|s-t|\frac{1}{q^{2}}-\frac{N^{2}}{qMZ_{y}}\right)$
$=M\left(N-\frac{(q-1)(q+1)}{3q}-\frac{N^{2}}{qMZ_{y}}\right).$ (32)
This completes the proof of (25a). Based on ${\theta}_{\text{max}}^{2}\left(M-\frac{1}{q}\right)\geq{\theta}_{c}^{2}(M-1)+{\theta}_{a}^{2}\left(1-\frac{1}{q}\right)$, we have (25b). ∎
A more optimal choice of $q$ can lead to further improvement of (31).
@@
${\theta}_{c}^{2}(M-1)+{\theta}_{a}^{2}\left(1-\frac{1}{\sqrt{\frac{3N^{2}}{MZ_{y}}}}\right)$ $\geq{\theta}_{c}^{2}(M-1)+{\theta}_{a}^{2}\left(1-\frac{1}{\hat{q}}\right)$
$\geq M\left(N-\frac{\hat{q}^{2}-1}{3q}-\frac{N^{2}}{\hat{q}MZ_{y}}\right)$
$\geq M\left(N-\frac{2N}{\sqrt{3MZ_{y}}}\right).$ (34)
@@ III-A2 Weight vector B ($\mathbf{w}^{B,q}$)
First, to facilitate the subsequent analysis, we consider a simplified version of the bound (25):
Corollary 4
For any weight vector $\mathbf{w}$ satisfying the weighting condition (14a), and any sequence set $(N,M,{\theta}_{\text{max}},\bm{\Gamma})\text{-}\mathcal{S}$,
${\theta}_{c}^{2}(M-1)+{\theta}_{a}^{2}\geq M\left(N-\mathcal{Q}\left(\mathbf{w},\frac{N^{2}}{MZ_{y}},\bm{0}_{Z_{x}}\right)\right),$ (35a)
${\theta}_{\text{max}}^{2}\geq N-\mathcal{Q}\left(\mathbf{w},\frac{N^{2}}{MZ_{y}},\bm{0}_{Z_{x}}\right).$ (35b)
Proof:
According to (25),
${\theta}_{\text{max}}^{2}$ $\geq\frac{{M}\left(N-\sum_{s,t=0}^{Z_{x}-1}l_{s,t,N}w_{s}w_{t}\right)-\frac{{N^{2}}}{Z_{y}}\sum_{i=0}^{Z_{x}-1}w_{i}^{2}}{M-\sum_{i=0}^{Z_{x}-1}w_{i}^{2}}$
$\geq\frac{{M}\left(N-\sum_{s,t=0}^{Z_{x}-1}l_{s,t,N}w_{s}w_{t}\right)-\frac{{N^{2}}}{Z_{y}}\sum_{i=0}^{Z_{x}-1}w_{i}^{2}}{M}$
$=N-\mathcal{Q}\left(\mathbf{w},\frac{N^{2}}{MZ_{y}},\bm{0}_{Z_{x}}\right).$ (36)
(35a) can be obtained similarly and hence omitted. ∎
Based on (35), a weight vector can be obtained by minimizing the following function using the Lagrange multiplier [31]:
$L_{q}(\mathbf{w})=\mathcal{Q}\left(\mathbf{w},\frac{N^{2}}{MZ_{y}},\bm{0}_{Z_{x}}\right)-2\lambda\left(\sum_{i=0}^{q-1}w_{i}-1\right),$ (37)
where $12\leq q\leq Z_{x}-$ 1, with $w_{i}=0$, for $i=q,q+1,\cdots,Z_{x}-1$. Let $\gamma=\arccos\left(1-\frac{MZ_{y}}{N^{2}}\right)$, $MZ_{y}\leq N^{2}$, and $q$ be an even positive integer such that $q{\gamma}\leq\pi+\gamma$. By setting $\gamma_{0}=\frac{\pi-q\gamma+\gamma}{2}$, one can obtain the following weight vector [^2] Although it looks similar to that in [13, Lemma 2] for aperiodic correlation, it is more generic as it considers the Doppler components in the aperiodic AF ($Z_{y}>1$). by relating the quadratic minimization solution of $L_{q}(\mathbf{w})$ to the Chebyshev polynomials of the second kind:
$w_{i}^{B,q}=\left\{\begin{array}[]{ll}\frac{\sin\frac{\gamma}{2}}{\sin\frac{q\gamma}{2}}\sin\left(\gamma_{0}+i\gamma\right),&0\leq i\leq q-1;\\ 0,&q\leq i\leq Z_{x}-1.\end{array}\right.$ (40)
Substituting $\mathbf{w}^{B,q}$ and $D=0$ into (25), the AF lower bounds below can be obtained.
Corollary 5
For any LAZ $\bm{\Gamma}$ such that $MZ_{y}\leq N^{2}$, and any integer $q$ satisfying $1\leq q<\text{min}\left\{Z_{x}+1,\frac{\pi}{\gamma}+1\right\}$, we have
${\theta}_{\text{max}}^{2}$ $\geq\frac{{\theta}_{c}^{2}(M-1)+{\theta}_{a}^{2}}{M}$
$\geq N-\frac{q-1}{2}-\frac{\sin\frac{q\gamma}{2}-\sin\frac{(q-2)\gamma}{2}}{2(1-\cos\gamma)\sin\frac{q\gamma}{2}}.$ (41)
Particularly, when $Z_{x}>\frac{\pi}{\gamma}$, one can minimize $L_{q}(\mathbf{w})$ in (37) over different values of $q$ by setting $q=\left\lfloor\frac{\pi}{\gamma}\right\rfloor+1$.
@@ III-B LAZ with $Z_{x}=N$ Case
Following the same steps as in (III-A) to (25), generalized aperiodic AF lower bounds for the LAZ $\bm{\Gamma}$ with $Z_{x}=N$ can be derived as follows:
@@
${\theta}_{c}^{2}(M-1)\left(1-\mathbf{w}^{T}\sum_{d=1}^{D}\bm{J}^{d}_{2N-1}\mathbf{w}\right)+{\theta}_{a}^{2}\left(1-\mathbf{w}^{T}\left(\bm{I}_{2N-1}+\sum_{d=1}^{D}\bm{J}^{d}_{2N-1}\right)\mathbf{w}\right)$
@@
  • When $Z_{y}=1$, the proposed bound (44b) reduces to the Arlery-Tan-Rabaste bound [15, Theorem 2] for aperiodic correlation, which exactly represents the generalized Levenshtein bound when $D=0$ [13, Theorem 1]. Again, these correlation bounds are equivalent to the AF bounds for $Z_{y}>1$ obtained using the non-optimal weight vector $\mathbf{p}=[1,0,\cdots,0]^{T}$, so the proposed bound (44b) is strictly tighter than them for $Z_{y}>1$.
@@ III-B1 Weight vector C ($\mathbf{w}^{C}$)
The proof of (46) is similar to that of Corollary 3 by substituting $\mathbf{w}^{C}$ and $D=0$ into (44b), note that $\sum_{s,t=0}^{2N-2}l_{s,t,N}w_{s}w_{t}=\frac{N(N-1)}{2N-1}$ for $\mathbf{w}^{C}$. ∎
@@
This confirms that the increased weight vector flexibility introduced by the new weighting condition (43) potentially allows the bound (44b) to be tighter than the bound (25b). When the conditions specified in Remark 6 are not met, there exist weight vectors superior to $\mathbf{w}^{C}$. The bound in (46b) can be improved by utilizing weight vectors like $\mathbf{w}^{A,q}$ and $\mathbf{w}^{B,q}$, but with a length of $12N-$ 1.
@@
$\exists\,\,(w_{s},w_{t})\,\,\text{s.t.}\sum_{\begin{subarray}{c}s,t=0\\ l_{s,t,N}=N-d\end{subarray}}^{2N-2}w_{s}w_{t}\neq 0,\quad{\theta}_{\text{max}}^{2}>{d^{2}},$ (47)
@@ IV Asymptotically Order-Optimal Sequence Set for AF in the LAZ
According to Corollary 1-7, $\theta_{\text{max}}$ should be at least $\mathcal{O}(\sqrt{N})$, when the order of the size of the LAZ (i.e. the order of $Z_{x}Z_{y}$) is no less than $\mathcal{O}(N)$ (see Table I for more examples). Therefore, a sequence set with $\theta_{\text{max}}$ asymptotically equal to $\mathcal{O}(\sqrt{N})$ is said to be “order-optimal” in this paper.
@@
Since $A_{\mathbf{s}}(\tau,\nu)={A_{\mathbf{s}}(-\tau,-\nu)^{\ast}}$, we only need to consider $\tau>0$ to analyze the maximum AF magnitude in a given LAZ. Note that if $\nu-a\tau\neq uN$ for any integer $u$, $\sum_{n=0}^{N-1}e^{\frac{j2\pi(\nu-a\tau)n}{N}}=0$ [28]. For $\tau>0$, we have
${\left|A_{\mathbf{s}}(\tau,\nu)\right|}^{2}$ $={\left|\sum_{n=0}^{N-1-\tau}{s}_{n}^{a}{\left({s}_{n+\tau}^{a}\right)}^{\ast}e^{\frac{j2\pi\nu n}{N}}\right|}^{2}=\left|\sum_{n=0}^{N-1-\tau}e^{\frac{j\pi an^{2}}{N}}e^{\frac{-j\pi a(n+\tau)^{2}}{N}}e^{\frac{j2\pi\nu n}{N}}\right|^{2}$
$=\left|\sum_{n=0}^{N-1-\tau}e^{\frac{j2\pi(\nu-a\tau)n}{N}}\right|^{2}=\left|\sum_{l=0}^{N-1-(N-\tau)}e^{\frac{-j2\pi(\nu+a(N-\tau))l}{N}}\right|^{2}$
$=\left|A_{\mathbf{s}}(N-\tau,-\nu)\right|^{2}.$ (50)
[…]
Case 1: $1\leq\tau\leq 0.45\sqrt{\frac{N}{|a|}}.$
According to Lemma 2 and for $\tau>0$, we have ${\left|A_{\mathbf{s}}(\tau,\nu)\right|}^{2}\leq(N-\tau)^{2}$. So in this case, we consider the range of $\tau$ satisfying ${\left|A_{\mathbf{s}}(\tau,\nu)\right|}^{2}\leq(N-\tau)^{2}\leq\frac{0.81N}{4|a|}$, which implies $N-0.45\sqrt{\frac{N}{|a|}}\leq\tau\leq N-1$. From the choice of $a$, $\nu$ and $\tau$, it is clear that $\nu-a\tau\neq uN$. According to (IV), we have
$\frac{{\left|A_{\mathbf{s}}(\tau,\nu)\right|}^{2}}{N}\leq\frac{0.2025}{|a|},\,\,\text{for}\;1\leq\tau\leq 0.45\sqrt{\frac{N}{|a|}}.$ (51)
Case 2: $10.45\sqrt{\frac{N}{|a|}}<\tau\leq\frac{\beta N}{|a|}-$ 1.
In this case, $0<-|a|+1+0.45\sqrt{|a|N}<\nu-a\tau<\beta N$. We introduce the condition $\sqrt{N}>\frac{|a|-1}{0.45\sqrt{|a|}}$, or simply, $N\geq 5|a|$. Note that
${\left|A_{\mathbf{s}}(\tau,\nu)\right|}^{2}$ $=\left|\sum_{n=0}^{N-1-\tau}e^{\frac{j2\pi(\nu-a\tau)n}{N}}\right|^{2}$
$=\left|\frac{1-e^{\frac{j2\pi(\nu-a\tau)(N-\tau)}{N}}}{1-e^{\frac{j2\pi(\nu-a\tau)}{N}}}\right|^{2}$
$=\frac{1-\cos\frac{2\pi(\nu-a\tau)(N-\tau)}{N}}{1-\cos\frac{2\pi(\nu-a\tau)}{N}}$
$=\frac{1-\cos\frac{2\pi(\nu-a\tau)\tau}{N}}{1-\cos\frac{2\pi(\nu-a\tau)}{N}}.$ (52)
Then, we divide Case 2 into the following two sub-cases:
Sub-case 2-1: $10.45\sqrt{\frac{N}{|a|}}<\tau\leq\frac{\beta N}{|a|}-$ 1 with $(\nu-a\tau)^{2}\geq\frac{5\beta^{2}|a|}{\sin^{2}(\beta\pi)}N$.
Let $\varphi=\frac{2\pi(\nu-a\tau)}{N}$, since $0<\nu-a\tau<\beta N$, we have $0<\varphi<2\pi\beta$. Note that $\sin\frac{\varphi}{2}\geq\frac{\sin\beta\pi}{2\beta\pi}\varphi$, so $1-\cos\frac{2\pi(\nu-a\tau)}{N}\geq 2\left(\frac{\sin\beta\pi}{2\beta\pi}\cdot\frac{2\pi(\nu-a\tau)}{N}\right)^{2}=\frac{2\sin^{2}(\beta\pi)}{\beta^{2}N}\cdot\frac{(\nu-a\tau)^{2}}{N}$. Then based on (IV), we have
$\frac{{\left|A_{\mathbf{s}}(\tau,\nu)\right|}^{2}}{N}$ $\leq\frac{2}{\frac{2\sin^{2}(\beta\pi)}{\beta^{2}N}\cdot\frac{(\nu-a\tau)^{2}}{N}N}=\frac{\beta^{2}}{\sin^{2}(\beta\pi)}\frac{N}{(\nu-a\tau)^{2}}$
$\leq\frac{\beta^{2}}{\sin^{2}(\beta\pi)}\frac{\sin^{2}(\beta\pi)}{5\beta^{2}|a|}=\frac{0.2}{|a|},$
for $\text{for}\;0.45\sqrt{\frac{N}{|a|}}$ $<\tau\leq\frac{\beta N}{|a|}-1\;\text{with}\;(\nu-a\tau)^{2}\geq\frac{5\beta^{2}|a|}{\sin^{2}(\beta\pi)}N.$ (53)
Sub-case 2-2: $10.45\sqrt{\frac{N}{|a|}}<\tau\leq\frac{\beta N}{|a|}-$ 1 with $(\nu-a\tau)^{2}<\frac{5\beta^{2}|a|}{\sin^{2}(\beta\pi)}N$.
In this sub-case, $\frac{\nu-a\tau}{N}\rightarrow 0$ as $N\rightarrow+\infty$. Following (IV), for $N\rightarrow+\infty$, we have
$=\frac{N\tau}{\pi(\nu-a\tau)}F\left(\frac{2\pi(\nu-a\tau)\tau}{N}\right),$
for $\text{for}\;0.45\sqrt{\frac{N}{|a|}}$ $<\tau\leq\frac{\beta N}{|a|}-1\;\text{with}\;(\nu-a\tau)^{2}<\frac{5\beta^{2}|a|}{\sin^{2}(\beta\pi)}N,$ (54)
where $F(\varphi)=\frac{1-\cos\varphi}{\varphi}$ with $\varphi=\frac{2\pi(\nu-a\tau)\tau}{N}=\frac{\tau}{\nu-a\tau}\frac{2\pi(\nu-a\tau)^{2}}{N}<\frac{2\pi\tau}{\nu-a\tau}\cdot\frac{5\beta^{2}|a|}{\sin^{2}(\beta\pi)}$. For $\varphi>2\pi$, $F(\varphi)\leq\frac{2}{\varphi}\leq\frac{1}{\pi}\approx 0.3183$. For $0<\varphi\leq 2\pi$, the maximum value of $F(\varphi)$ equals $F(\varphi_{0})\approx 0.7246$, $\varphi_{0}\approx 2.3311$ [23].
Then, we consider the $\frac{\tau}{\nu-a\tau}$ part in (IV). Note that $10.45\sqrt{\frac{N}{|a|}}<\tau\leq\frac{\beta N}{|a|}-$ 1, we have $\frac{1}{\frac{-a-1}{0.45\sqrt{N/|a|}}-a}\leq\sqrt{\frac{N}{|a|}}\leq\frac{1}{\frac{a+1}{\beta N/|a|-1}-a}$. For $N\rightarrow+\infty$, we have $\frac{1}{\frac{-a-1}{0.45\sqrt{N/|a|}}-a}\rightarrow\frac{1}{|a|}$ and $\frac{1}{\frac{a+1}{\beta N/|a|-1}-a}\rightarrow\frac{1}{|a|}$, so $\frac{\tau}{\nu-a\tau}\rightarrow\frac{1}{|a|}$. Thus, from (IV), we have
$\lim_{N\to+\infty}\frac{\max_{(\tau,\nu)}{\left|A_{\mathbf{s}}(\tau,\nu)\right|^{2}}}{N}=\frac{F(\varphi_{0})}{\pi|a|}\approx\frac{0.2306}{|a|},$
for $\text{for}\;0.45\sqrt{\frac{N}{|a|}}<\tau\leq\frac{\beta N}{|a|}-1\;\text{with}\;(\nu-a\tau)^{2}<\frac{5\beta^{2}|a|}{\sin^{2}(\beta\pi)}N.$ (55)
Overall, based on the results from (51), (IV), and (IV), we conclude (49). ∎
@@
We then analyze the maximum CAF magnitude in the LAZ for Chu sequences. Consider two Chu sequence $\mathbf{s}^{1}=\left[{s}^{a_{1}}_{0},{s}^{a_{1}}_{1},\cdots,{s}^{a_{1}}_{N-1}\right]$ and $\mathbf{s}^{2}=\left[{s}^{a_{2}}_{0},{s}^{a_{2}}_{1},\cdots,{s}^{a_{2}}_{N-1}\right]$ with $a_{1}>a_{2}$. We need the following lemma based on the Van Der Corput method [33].
Lemma 5
@@
${\left|A_{\mathbf{s}^{1},\mathbf{s}^{2}}(\tau,\nu)\right|}\leq 3\left(\sqrt{a_{1}-a_{2}}+\frac{2}{\sqrt{a_{1}-a_{2}}}\right)\sqrt{N}-\frac{3|\tau|\sqrt{a_{1}-a_{2}}}{\sqrt{N}}.$ (57)
@@
${\left|A_{\mathbf{s}^{1},\mathbf{s}^{2}}(\tau,\nu)\right|}^{2}$ $=\left|\sum_{n=0}^{N-1-\tau}s^{1}_{n}s_{n+\tau}^{2\ast}e^{\frac{j2\pi\nu n}{N}}\right|^{2}=\left|\sum_{n=0}^{N-1-\tau}e^{j\pi\frac{(a_{1}-a_{2})n^{2}+2(\nu-a_{2}\tau)n-a_{2}\tau^{2}}{N}}\right|^{2}$
$=\left|\sum_{n=0}^{N-1-\tau}e^{j\pi\frac{(a_{1}-a_{2})n^{2}+2(\nu-a_{2}\tau)n}{N}}\right|^{2}.$ (58)
Similarly, we have ${\left|A_{\mathbf{s}^{1},\mathbf{s}^{2}}(\tau,\nu)\right|}^{2}=\left|\sum_{n=0}^{N-1-\tau}e^{j\pi\frac{(a_{1}-a_{2})n^{2}+2(\nu-a_{1}\tau)n}{N}}\right|^{2}$, for $\tau<0$.

### v3

@@ III-A1 Weight vector A ($\mathbf{w}^{A,q}$)
By appropriately selecting $q$, we obtain:
@@
See Appendix A. ∎
@@ III-A2 Weight vector B ($\mathbf{w}^{B,q}$)
We derive the weight vector $\mathbf{w}^{B,q}$ using a quadratic minimization approach, similar to [13, Lemma 2] and [31]. The resulting weight vector is expressed as:
$w_{i}^{B,q}=\left\{\begin{array}[]{ll}\frac{\sin\frac{\gamma}{2}}{\sin\frac{q\gamma}{2}}\sin\left(\gamma_{0}+i\gamma\right),&0\leq i\leq q-1;\\ 0,&q\leq i\leq Z_{x}-1,\end{array}\right.$ (32)
where $q$ is an even positive integer satisfying $q\gamma\leq\pi+\gamma$, and $\gamma_{0}=\frac{\pi-q\gamma+\gamma}{2}$. Here, $\gamma=\arccos\left(1-\frac{MZ_{y}}{N^{2}}\right)$ and $MZ_{y}\leq N^{2}$.
By substituting $\mathbf{w}^{B,q}$ and $D=0$ into the bound (22), and selecting $q=\left\lfloor\frac{\pi}{\gamma}\right\rfloor+1$, we obtain the following result.
@@ III-B LAZ with $Z_{x}=N$ Case
Following the same steps as in (III-A) to (22), replacing $l_{s,t,N}=|t-s|$ with the extended definition $l^{\prime}_{s,t,N}=\min\left\{|t-s|,2N-1-|t-s|\right\}$, generalized aperiodic AF lower bounds for the LAZ $\bm{\Gamma}$ with $Z_{x}=N$ can be derived as follows:
@@
${\theta}_{c}^{2}(M-1)\left(1-\mathbf{w}^{T}\sum_{d=1}^{D}\bm{J}^{d}_{2N-1}\mathbf{w}\right)$
$+$ ${\theta}_{a}^{2}\left(1-\mathbf{w}^{T}\left(\bm{I}_{2N-1}+\sum_{d=1}^{D}\bm{J}^{d}_{2N-1}\right)\mathbf{w}\right)$
@@
  • When $Z_{y}=1$, the proposed bound (35b) reduces to the Arlery-Tan-Rabaste bound [15, Theorem 2], which corresponds to the Levenshtein bound when $D=0$ [13, Theorem 1]. Again, these correlation bounds are equivalent to the AF bounds for $Z_{y}>1$ obtained using the non-optimal weight vector $\mathbf{p}=[1,0,\cdots,0]^{T}$, so the proposed bound (35b) is strictly tighter than them for $Z_{y}>1$.
@@ III-B1 Weight vector C ($\mathbf{w}^{C}$)
The proof of (37) is similar to that of Corollary 2 by substituting $\mathbf{w}^{C}$ and $D=0$ into (35b), note that $\sum_{s,t=0}^{2N-2}l^{\prime}_{s,t,N}w_{s}w_{t}=\frac{N(N-1)}{2N-1}$ for $\mathbf{w}^{C}$. ∎
@@
$\exists\,\,(w_{s},w_{t})\,\,\text{s.t.}\sum_{\begin{subarray}{c}s,t=0\\ l^{\prime}_{s,t,N}=N-d\end{subarray}}^{2N-2}w_{s}w_{t}\neq 0,\quad{\theta}_{\text{max}}^{2}>{d^{2}},$ (38)
@@ IV Asymptotically Order-Optimal Sequence Set for AF in the LAZ
According to Corollary 1-4, $\theta_{\text{max}}$ should be at least $\mathcal{O}(\sqrt{N})$, when the order of the size of the LAZ (i.e. the order of $Z_{x}Z_{y}$) is no less than $\mathcal{O}(N)$ (see Table I for more examples). Therefore, a sequence set with $\theta_{\text{max}}$ asymptotically equal to $\mathcal{O}(\sqrt{N})$ is said to be “order-optimal” in this paper.
@@
See Appendix B. ∎
@@
We then analyze the maximum CAF magnitude in the LAZ for Chu sequences. Consider two Chu sequence $\mathbf{s}^{1}=\left[{s}^{a_{1}}_{0},{s}^{a_{1}}_{1},\cdots,{s}^{a_{1}}_{N-1}\right]$ and $\mathbf{s}^{2}=\left[{s}^{a_{2}}_{0},{s}^{a_{2}}_{1},\cdots,{s}^{a_{2}}_{N-1}\right]$ with $a_{1}>a_{2}$. We need the following lemma on the bounding method of Van Der Corput [33].
Lemma 5 ([33, Theorem 8.15])
@@
${\left|A_{\mathbf{s}^{1},\mathbf{s}^{2}}(\tau,\nu)\right|}$
$\leq$ $3\left(\sqrt{a_{1}-a_{2}}+\frac{2}{\sqrt{a_{1}-a_{2}}}\right)\sqrt{N}-\frac{3|\tau|\sqrt{a_{1}-a_{2}}}{\sqrt{N}}.$ (42)
@@
${\left|A_{\mathbf{s}^{1},\mathbf{s}^{2}}(\tau,\nu)\right|}$ $=\left|\sum_{n=0}^{N-1-\tau}s^{1}_{n}s_{n+\tau}^{2\ast}e^{\frac{j2\pi\nu n}{N}}\right|$
$=\left|\sum_{n=0}^{N-1-\tau}e^{j\pi\frac{(a_{1}-a_{2})n^{2}+2(\nu-a_{2}\tau)n-a_{2}\tau^{2}}{N}}\right|$
$=\left|\sum_{n=0}^{N-1-\tau}e^{j\pi\frac{(a_{1}-a_{2})n^{2}+2(\nu-a_{2}\tau)n}{N}}\right|.$ (43)
Similarly, we have ${\left|A_{\mathbf{s}^{1},\mathbf{s}^{2}}(\tau,\nu)\right|}=\left|\sum_{n=0}^{N-1-\tau}e^{j\pi\frac{(a_{1}-a_{2})n^{2}+2(\nu-a_{1}\tau)n}{N}}\right|$, for $\tau<0$.